# Disjoint supercyclicity in Banach algebras

### v7

@@ 3.1. Applications to generalized weighted shifts on Hilbert C*-modules
The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\in\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
@@
By [21, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
@@
Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ We will assume that for every compact subset $K$ of $X$ and every $S\in\mathcal{F}$ there exists some $F\in\mathcal{F}$ with $F\subseteq S$ such that
$K\cap\alpha_{n,\ell}(K)=\varnothing\quad\text{and}\quad\alpha_{n,\ell}^{-1}(\alpha_{n,s}(K))\cap K=\varnothing$
for all $n\in F$ and each distinct $s,\ell\in\{1,\dots,N\}.$ Also, we consider families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
@@
By some calculations similar to those in the proof of [16, Theorem 3.2] and [17, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 2.2 are satisfied. ∎
@@
3.3. Applications to generalized weighted shifts on function sequence spaces
In this subsection, we assume that $\Omega=\mathbb{R}.$ and we let now $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [15], for $\tau\in C_{b}(\mathbb{R}),$ we put
$\mathcal{C}_{\tau}:=\{f\in\mathcal{C}:\,\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}<\infty\}.$
For each $f\in\mathcal{C}_{\tau}$ we define
$\|f\|_{\tau}:=\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}.$
Then, $\mathcal{C}_{\tau}$ is a Banach algebra [14]. We will call this algebra Segal algebra corresponding to $\tau$.
As observed in [15], for a homeomorphism $\alpha$ of $\mathbb{R},$ we consider corresponding Segal algebra $\mathcal{C}_{\tau\circ\alpha^{k}}$ for all $k\in\mathbb{Z}$. Clearly, $f\cdot(\tau\circ\alpha^{k})\in\mathcal{C}$ for all $k\in\mathbb{Z}$ whenever $f\in C_{0}(\mathbb{R})$ since $\tau\circ\alpha^{k}\in C_{b}(\mathbb{R})$ for all $k\in\mathbb{Z}.$ By definition, $\|f\|_{\infty}\leq\|f\|_{\tau\circ\alpha^{k}}$ whenever $f\in\mathcal{C}_{\tau\circ\alpha^{k}}.$ We recall now the following lemmas from [15].
@@ 3.3. Applications to generalized weighted shifts on function sequence spaces
[15, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
@@
    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}=1$ on $K_{\epsilon_{2}}^{(\tau)}$,

### v8

@@ 4. Applications
The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
@@
By [29, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
@@
Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ and families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
@@
By some calculations similar to those in the proof of [20, Theorem 3.2] and [21, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 3.2 are satisfied. ∎
@@
5. Dynamics on weighted spaces of Radon measures
If $T_{\alpha,w}$ denotes the weighted composition operator on $C_{0}(\Omega)$ with respect to $\alpha$ and $w,$ that is $T_{\alpha,w}(f)=w\cdot(f\circ\alpha)$ for all $f\in C_{0}(\Omega),$ then the adjoint $T_{\alpha,w}^{*}$ is a bounded operator on $M(\Omega)$ where $M(\Omega)$ stands for the Banach space of all Radon measures on $\Omega$ equipped with the total variation norm. It is straightforward to check that
$T_{\alpha,w}^{*}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ Here $\mu\ \circ\alpha^{-1}(E)=\mu(\alpha^{-1}(E))$ for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ By $(1)$ and $(2)$ it follows then that for every $n\in\mathbb{N},$ $\mu\in M(\Omega)$ and a Borel measurable subset $E\subseteq\Omega$ we have that
${T_{\alpha,w}^{*n}}(\mu)(E)=\int_{E}\prod_{j=0}^{n-1}w\circ\alpha^{j-n}\,d\mu\ \circ\alpha^{-n}$
and
${S_{\alpha,w}^{*n}}(\mu)(E)=\int_{E}\prod_{j=1}^{n}(w\circ\alpha^{n-j})^{-1}\,d\mu\ \circ\alpha^{n}.$
Throughout this paper, $M_{+}(\Omega)$ will denote the set of all non-negative Radon measures on $\Omega.$ For every Radon measure $\mu$ on $\Omega,$ we let as usual $|\mu|$ denote the total variation of $\mu.$ A weight on $\Omega$ with respect to $\alpha$ is a continuous positive function $\rho$ on $\Omega$ satisfying that
$(\rho\circ\alpha)(t)\leq C\rho(t)\quad\text{and}\quad(\rho\circ\alpha^{-1})(t)\leq C\rho(t)$
for all $t\in\Omega$ and some constant $C>0$.
If $\mu\in M(\Omega)$ and $\rho\in L^{1}(|\mu|)$, we let $\mu^{(\rho)}$ be the Borel measure on $\Omega$ given by
$\mu^{(\rho)}(E)=\int_{E}\rho\,d\mu$
for every Borel subset $E$ of $\Omega$ and we set
$M_{\rho}(\Omega):=\left\{\mu\in M(\Omega)\;\middle|\;\rho\in L^{1}(|\mu|)\text{ and }\mu^{(\rho)}\text{ is regular}\right\}.$
Let $\|\cdot\|_{\rho}:M_{\rho}(\Omega)\to\mathbb{R}^{+}$ be given by
$\|\mu\|_{\rho}=|\mu^{(\rho)}|(\Omega).$
@@ 5. Dynamics on weighted spaces of Radon measures
$M_{\rho}(\Omega)$ is a subspace and $\|\cdot\|_{\rho}$ is a norm on $M_{\rho}(\Omega).$
Proof.
Let $\mu,\nu\in M_{\rho}(\Omega)$ and $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$ be the Jordan decomposition of $\mu.$ It is easy to check that $|\mu+\nu|\leq|\mu|+|\nu|.$ Now, since $\rho\in L^{1}(|\mu|)\cap L^{1}(|\nu|),$ we get that $\rho\in L^{1}(|\mu+\nu|).$ Next, we have that $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i\big(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)}\big),$ is the Jordan decomposition of $\mu^{(\rho)}.$
Therefore, $\mu_{1}^{(\rho)}+\mu_{2}^{(\rho)},$ $\mu_{3}^{(\rho)}+\mu_{4}^{(\rho)}\leq|\mu^{(\rho)}|,$ hence, if $|\mu^{(\rho)}|=0,$ then $\mu_{j}^{(\rho)}=0$ for all $j\in\{1,\ldots,4\},$ but this means that $\mu_{j}=0$ for all $j\in\{1,\ldots,4\}$ since $\rho$ is a positive continuous function and $\mu_{j}$ is a Borell measure. Thus, $\mu=0$ in this case. ∎
Lemma 5.2.
Let $K$ be a compact subset of $\Omega,$ $\delta>0$ and $\mu\in M_{\rho}(\Omega).$ If $|\mu^{(\rho)}|(K)<\delta,$ then $|\mu|(K)\leq\dfrac{2\delta}{m_{K}},$ where $m_{K}:=\inf_{x\in K}\rho(x)>0.$
Proof.
Let again $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$ be the Jordan decomposition of $\mu,$ then $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)})$ is the Jordan decomposition of $\mu^{(\rho)}.$ If $|\mu^{(\rho)}|(K)<\delta,$ then $(\mu_{1}^{(\rho)}+\mu_{2}^{(\rho)})(K)<\delta$ and $(\mu_{3}^{(\rho)}+\mu_{4}^{(\rho)})(K)<\delta.$ Since $\rho$ is positive and continuous, we must have that $m_{K}>0,$ hence, the above inequalities give $(\mu_{1}+\mu_{2})(K)<\dfrac{\delta}{m_{K}}$ and $(\mu_{3}+\mu_{4})(K)<\dfrac{\delta}{m_{K}}.$ Now, since $|\mu|\leq\sum_{j=1}^{4}\mu_{j},$ the statement follows. ∎
Remark 5.3.
If $\mu\in M_{\rho}(\Omega)$ and $\delta>0,$ it is not hard to check that there exists a compact subset $K\subseteq\Omega$ and $\nu\in M(\Omega)$ such that
$|\nu|(K^{c})=0\quad\text{and}\quad|\mu^{(\rho)}-\nu^{(\rho)}|(\Omega)<\delta.$
Indeed, since $\mu^{(\rho)}$ is a regular Borel measure, there exists a compact subset $K\subseteq\Omega$ such that
$|\mu^{(\rho)}|(K^{c})<\frac{\delta}{2},$
hence we can let $\nu$ be the Borel measure on $\Omega$ given by
$\nu(E):=\mu(E\cap K),$
for every Borel subset $E\subseteq\Omega.$
Lemma 5.4.
The operators $T_{\alpha,w}^{*}$ and $S_{\alpha,w}^{*}$ are bounded linear self-mappings on $M_{\rho}(\Omega).$
Proof.
Let $\mu\in M_{\rho}(\Omega)$ and
$\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$
be the Jordan decomposition of $\mu.$ Then $\rho\in\bigcap_{j=1}^{4}L^{1}(\mu_{j}).$ By the change of variables, we obtain
$\int\rho\,(w\circ\alpha^{-1})\,d(\mu_{j}\circ\alpha^{-1})=\int(\rho\circ\alpha)\,w\,d\mu_{j}\leq$
$C\Big(\sup_{t\in\Omega}w(t)\Big)\int\rho\,d\mu_{j}<\infty,$
for all $j\in\{1,2,3,4\},$ because
$(\rho\circ\alpha^{-1})(t)\leq C\rho(t),\quad\text{for all }t\in\Omega,$
by the assumption. Thus, $\rho\in L^{1}\big(T_{\alpha,w}^{*}(\mu_{j})\big)$ for all $j\in\{1,2,3,4\}.$ Since
$T^{*}_{\alpha,w}(\mu)=T^{*}_{\alpha,w}(\mu_{1})-T^{*}_{\alpha,w}(\mu_{2})+i\Big(T^{*}_{\alpha,w}(\mu_{3})-T^{*}_{\alpha,w}(\mu_{4})\Big)$
is the Jordan decomposition of $T^{*}_{\alpha,w}(\mu)$ by the arguments from the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, it follows that $\rho\in L^{1}\big(|T_{\alpha,w}^{*}(\mu)|\big).$
Next, given $\epsilon>0$ and a Borel subset $E$ of $\Omega,$ choose some open subset $\mathcal{O}\subseteq E$ and a compact subset $K\subseteq\Omega$ such that $K$ and $\alpha^{-1}(E)$ are subsets of $\mathcal{O}$ and
$|\mu^{(\rho)}|(\mathcal{O}\setminus\alpha^{-1}(E)),|\mu^{(\rho)}|(\mathcal{O}\setminus K)<\varepsilon.$
Since $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i\big(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)}\big)$ is the Jordan decomposition of $\mu^{(\rho)}$, we have that $|\mu_{j}^{(\rho)}|\leq|\mu^{(\rho)}|$ for all $j\in\{1,\dots,4\}$, hence
$|\mu_{j}^{(\rho)}|\big(\mathcal{O}\setminus\alpha^{-1}(E)\big)<\varepsilon$
and
$|\mu_{j}^{(\rho)}|\big(\mathcal{O}\setminus K\big)<\varepsilon$
for all $j\in\{1,\dots,4\}$. By the previous inequality, we get that
$\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)>C\Big(\sup_{t\in\Omega}w(t)\Big)\int_{\mathcal{O}\setminus\alpha^{-1}(E)}\rho\,d\mu_{j}\geq\int_{\alpha(\mathcal{O})\setminus(E)}\rho\big(w\circ\alpha^{-1}\big)\,d\big(\mu_{j}\circ\alpha^{-1}\big),$
and similarly,
$\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)\geq\int_{\alpha(\mathcal{O})\setminus\alpha(K)}\rho\big(w\circ\alpha^{-1}\big)\,d\big(\mu_{j}\circ\alpha^{-1}\big)$
for all $j\in\{1,\dots,4\}$.
Since
$\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}=\big(T^{*}_{\alpha,w}(\mu_{1})\big)^{(\rho)}-\big(T^{*}_{\alpha,w}(\mu_{2})\big)^{(\rho)}+i\Big(\big(T^{*}_{\alpha,w}(\mu_{3})\big)^{(\rho)}-\big(T^{*}_{\alpha,w}(\mu_{4})\big)^{(\rho)}\Big)$
is the Jordan decomposition of $\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}$, we deduce that
$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\big(\alpha(\mathcal{O})\setminus E)\big)<4\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)$
and
$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\big(\alpha(\mathcal{O})\setminus\alpha(K)\big)<4\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big).$
Now, $\alpha(\mathcal{O})$ is open and $\alpha(K)$ is compact because $\alpha$ is a homeomorphism, and, moreover, $E\subseteq\alpha(\mathcal{O}),$ hence $\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}$ is regular.
Finally, by the above,
$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\leq 4C\Big(\sup_{t\in\Omega}w(t)\Big)|\mu^{(\rho)}|.$
Theorem 5.5.
@@
    The set of periodic elements of $T_{\alpha,\omega}^{*}$ is dense in $M_{\rho}(\Omega)$.